OpenStudy (anonymous):

60 small fish =5 big fish 40 big fish =4 fishing nets 4 fishing nets= 1 raft how many small fish = 1 raft?

OpenStudy (welshfella):

working backwards from the last line 1 raft = 4 fishing nets = 40 big fish 40 big fish is equal; to how many small fish?

OpenStudy (anonymous):

400

OpenStudy (welshfella):

nope 5 big fish = 60 small fish 1 big fish = 60/5 = 12 small fish so 40 big fish = 40 * 12 = 480 small fish

OpenStudy (anonymous):

shoot! I knew that duh ok thanks

OpenStudy (welshfella):

lol yw